A football coach claims that the average weight of all the opposing teams’ members is 225 pounds. For a test of the claim, a sample of 50 players is taken from all the opposing teams. The mean is found to be 230 pounds, and the standard deviation is 15 pounds. At α = 0.01, test the coach’s claim. Solve using the traditional method. Also calculate the p-value for your answer.

             = 230

                        S= 15

                        n= 50
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :   = 225

Ha :    225

= 230

= 225

s = 15

n = 50

degrees of freedom = n - 1 = 50 - 1 =49

Test statistic = t

= ( - ) / s / n

= (230 - 225) / 15 / 50

= 2.36

Test statistic = 2.36

P-value = 0.0223

= 0.01

P-value >  

Fail to reject the null hypothesis .

The coach’s claim is not true .